# FAST-Splat: Fast, Ambiguity-Free Semantics Transfer in Gaussian Splatting

### v1

@@ Abstract
We present FAST-Splat for fast, ambiguity-free semantic Gaussian Splatting, which seeks to address the main limitations of existing semantic Gaussian Splatting methods, namely: slow training and rendering speeds; high memory usage; and ambiguous semantic object localization. In deriving FAST-Splat, we formulate open-vocabulary semantic Gaussian Splatting as the problem of extending closed-set semantic distillation to the open-set (open-vocabulary) setting, enabling FAST-Splat to provide precise semantic object localization results, even when prompted with ambiguous user-provided natural-language queries. Further, by exploiting the explicit form of the Gaussian Splatting scene representation to the fullest extent, FAST-Splat retains the remarkable training and rendering speeds of Gaussian Splatting. Specifically, while existing semantic Gaussian Splatting methods distill semantics into a separate neural field or utilize neural models for dimensionality reduction, FAST-Splat directly augments each Gaussian with specific semantic codes, preserving the training, rendering, and memory-usage advantages of Gaussian Splatting over neural field methods. These Gaussian-specific semantic codes, together with a hash-table, enable semantic similarity to be measured with open-vocabulary user prompts and further enable FAST-Splat to respond with unambiguous semantic object labels and $3$D masks, unlike prior methods. In experiments, we demonstrate that FAST-Splat is $4$x to $6$x faster to train with a $13$x faster data pre-processing step, achieves between $18$x to $75$x faster rendering speeds, and requires about $3$x smaller GPU memory, compared to the best-competing semantic Gaussian Splatting methods. Further, FAST-Splat achieves relatively similar or better semantic segmentation performance compared to existing methods. After the review period, we will provide links to the project website and the codebase.
{shorinwa,jksun,schwager}@stanford.edu
@@ 1 Introduction
Research breakthroughs in vision-language foundation models have underpinned the remarkable performance of state-of-the-art object detection and classification [24, 50], segmentation [29, 15], and image captioning [27, 43]. In general, these models learn useful multi-modal image-language representations, entirely supervised with $2$D image-text pairs, all within a shared representation space. Recent work has shown that grounding the semantic knowledge encoded by vision-language foundation models in $3$D can be useful in improving semantic segmentation and localization [12, 35] (especially in models that take in low-resolution images as input), enabling applications such as $3$D scene-editing [14, 39] and robotic manipulation [28, 36]. Although these methods demonstrate compelling object segmentation performance from open-vocabulary queries, these methods fail to disambiguate between the semantic classes of unique objects that are semantically-similar to the natural language query [12]. For example, when a user queries for “tea,” existing semantic Gaussian Splatting methods localize the coffee machine without providing any clarifying information on the identity of the localized object, suggesting that “tea” exists in the scene, when the contrary is true. This ambiguity generally arises from the fact that many vision-language foundation methods, e.g., CLIP, utilize a bag-of-words approach in measuring the semantic similarity between image-language queries. Moreover, many semantic methods for Gaussian Splatting are notably slow, requiring a significant amount of memory and computation time for training and inference. We seek to derive a semantic Gaussian Splatting method that addresses these limitations.
@@
In this work, we propose FAST-Splat, a Fast, Ambiguity-free Semantics Transfer method for Gaussian Splatting. FAST-Splat generates fine-grained semantic localization, with highly specific semantic classes even in the presence of ambiguous natural-language prompts, and as its name implies, runs much faster than existing semantic Gaussian Splatting methods. FAST-Splat utilizes the vision-language foundation model CLIP [27] to extend closed-set semantic Gaussian Splatting to an open-world setting, while retaining the non-ambiguity of localized objects inherent in closed-set methods. Unlike existing semantic Gaussian Splatting methods which rely on neural-field models for semantic distillation, FAST-Splat directly augments each Gaussian with specific semantic codes, without any neural model. This key insight not only enables FAST-Splat to exploit the explicit scene-representation of Gaussian Splatting for faster training and rendering speeds, but also enables FAST-Splat to provide unambiguous semantic object labels in response to open-vocabulary user queries, e.g., in semantic object localization, upon identifying semantically-similar objects using the semantic codes associated with each Gaussian, along with the semantic hash-table.
@@
We compare FAST-Splat to existing semantic Gaussian Splatting methods [26, 49]. FAST-Splat achieves $4$x to $6$x faster training times and $18$x to $75$x faster rendering speeds, while requiring $3$x smaller GPU memory. Meanwhile, FAST-Splat achieves competitive semantic object localization performance, outperforming prior work in some scenes, while achieving the second-best performance in the other scenes. Notably, FAST-Splat enables semantic disambiguation in object localization, even with ambiguous user-provided natural-language queries, which we illustrate in LABEL:fig:teaser. In contrast to prior work, FAST-Splat provides a clarifying semantic label to each localized object, disambiguating the semantic identity of the localized object. For example, in LABEL:fig:teaser, when prompted with the query “tea,” FAST-Splat notifies a user that it localized a coffee machine, and an apple when given the query: “fruit.” Such disambiguation can be critical to enabling interesting downstream applications, e.g., in robotics.
@@
  • We introduce FAST-Splat, a Fast, Ambiguity-Free Semantic Gaussian Splatting method that enables notably faster ($4$x to $6$x faster) language-semantics grounding in $3$D scenes, distilled from $2$D vision models.
@@
  • FAST-Splat achieves superior rendering speeds for semantic Gaussian Splatting, about $18$x to $75$x faster rendering speeds compared to prior work.
@@ 2 Related Work
Grounding Language in 3D. To bridge the gap between $2$D object detection and $3$D object localization, prior work has examined grounding semantic embeddings from pre-trained image and text encoders in the $3$D world. CLIP-Fields [35] learns an implicit spatial mapping from $3$D points to CLIP image embeddings and Sentence-Bert [30] text embeddings using back-projected RGB-D points. VLMaps [10] and NLMap [3] distill vision-language features into grid-based maps. The sparsity of these maps limits the resolution of the open-vocabulary segmentation results achievable by these methods. For high-resolution object segmentation, Distilled Features Fields (DFF) [14], CLIP-NeRF [39], and LERF [12] ground CLIP [27], LSeg [15], and DINO [45] features in a NeRF representation of the scene. Moreover, similar techniques have been employed in [26, 9, 49, 51] to distill semantic features into Gaussian Splatting representations. The resulting distilled radiance fields (DRFs) enable dense open-vocabulary object segmentation in $3$D and have been applied to scene-editing [14, 39] and robotic manipulation [28, 36, 48, 37]. Despite their success in object segmentation, existing DRFs fail to precisely identify the semantic class of the segmented object, e.g., for a given query, these methods localize all objects related to the query without resolving the unique semantic identity of each object. This ambiguity creates a gap in scene understanding, which is often required in many downstream tasks, e.g., in robotic exploration or manipulation. In this work, we introduce a DRF with fine-grained semantics, enabling precise semantic identification of objects in a scene, disambiguating language-guided semantic object localization, and addressing existing challenges.
@@ 4 Fast, Ambiguity-Free Semantic Transfer
where ${\lambda\in[0,1]}$ denotes a weighting parameter between the ${\ell_{1}}$-loss term and the differentiable structural similarity index measure (D-SSIM) loss term, and $\mathcal{I}$ and $\hat{\mathcal{I}}$ denote the ground-truth and predicted RGB image, respectively. The Gaussian Splatting model is trained using the Adam optimizer [13], a gradient-based optimization method.
@@
Almost all existing semantic distillation methods rely on neural architectures for grounding language in $3$D in NeRFs [39, 14, 12] and Gaussian Splatting scenes [26, 9, 49, 51]. Although effective in enabling open-set semantic object localization, neural-based approaches introduce significant computational bottlenecks into the training pipeline of NeRF and Gaussian Splatting scenes. Moreover, when utilized in Gaussian Splatting, neural-based semantic distillation essentially erases the computational gains, e.g., faster training times and rendering rates, afforded by Gaussian Splatting when compared to NeRFs.
@@
Via FAST-Splat, we seek to derive a method for open-set semantic distillation, while preserving the faster training times and rendering rates achieved by the underlying Gaussian Splatting method. To achieve this goal, we revisit the fundamental insight exploited by the authors of the original Gaussian Splatting work [11], namely: that only representing occupied space explicitly can lead to significant speedup in training and rendering times without any notable degradation in the visual fidelity of the representation. Specifically, Gaussian Splatting harnesses the explicit form and sparsity of an ellipsoidal-primitive-based scene representation to achieve faster training times and rendering times compared to NeRFs, the state-of-the-art at the time, which relied on implicit fields for the color and density of the scene. By learning neural encoders for dimensionality reduction or neural fields for language semantics, existing semantic distillation methods for Gaussian Splatting fail to take advantage of this insight. In contrast, in FAST-Splat, we leverage this insight to speed up training and rendering times, by eschewing neural models. In particular, we augment the attributes of each ellipsoid with a semantic parameter, associated with its semantic identity ${\beta\in\mathbb{R}^{3}}$, which is trained simultaneously with the ellipsoid’s other attributes.
@@
Closed-Set Semantic Gaussian Splatting. We formulate the open-set semantic distillation problem as the problem of extending closed-set semantic distillation (segmentation) methods to the open-set regime. To obtain the closed-set of semantic categories, we feed each image ${I\in\mathcal{D}}$ into an object detector, e.g., YOLO [40], DETR [2], and Efficient DETR [42], to generate a list of objects present in the image. Although closed-set object detectors generally exhibit remarkable performance, these detectors occasionally fail to identify all instances of an object in an image, given that the size of their dictionary of detectable objects is often limited. To alleviate this issue, we further process the outputs of the object detector using an open-set object detector, e.g., GroundingDINO [22], with the input prompt given by the classes of objects identified by the object detector. In preliminary results, we find that this additional data pre-processing stage improves the robustness of the data pre-processing pipeline, by eliminating false detections made by the closed-set object detector.
Figure 1: Unlike prior semantic Gaussian Splatting methods, FAST-Splat trains a non-neural-based semantic Gaussian Splatting model to achieve much faster training and rendering times by extending closed-set semantic segmentation to the open-vocabulary setting. FAST-Splat distills language semantics from a closed-set object detector, e.g., YOLO, augmented with an open-set object detector, e.g., GroundingDINO, and a image segmentation model, e.g., SAM-2.
@@
At this stage, each image has a list of constituent objects and bounding boxes specifying the location of these objects in the image. However, high-fidelity semantic distillation generally requires pixel-wise semantic classes. To generate dense semantic maps, we utilize SAM-2 [29], passing in each object class and its associated bounding box. We maintain a dictionary (hash table), representing the closed-set of objects identified in the image. Now, we can distill the pixel-wise semantic information into the underlying Gaussian Splatting model. To predict the semantic class of each ellipsoid in the scene, we apply an affine transformation $\mathcal{W}$ to the semantic attribute $\beta$ of each ellipsoid, mapping $\beta$ to an $N$-dimensional space, where $N$ denotes the size of the dictionary. Subsequently, we apply the softmax function to the resulting outputs to define a probability over the entries in the dictionary. We extend this procedure to view-based rendering. Given a camera pose for rendering, we utilize differentiable tile-based rasterization [11] to render a semantic feature map corresponding to the camera pose. We utilize the aforementioned procedure to transform the semantic feature of each pixel to a probability distribution over the entries in the closed-set of object classes. We optimize the semantic attributes using the multi-class cross-entropy loss function, given by:
@@
where $\mathcal{T}$ represents the set of all indices of pixels in the rendered feature map $\mathcal{F}_{\mathcal{I}}$ associated with image $\mathcal{I}$ (after applying the affine transformation $\mathcal{W}$), $c$ denotes the true class of the pixel, and $\gamma_{p}$ represents the weight associated with pixel $p$. We augment the set of object classes for each view with an undetected class label and assign a lower weight to pixels with this class label, i.e., pixels that were not segmented during the data pre-processing phase. This augmentation step serves as a regularization term for the stability of undetected pixels. We train all attributes of the Gaussian Splatting representation, e.g., the geometric, visual, and semantic attributes and the parameters in $\mathcal{W}$, simultaneously using the loss function: ${\mathcal{L}_{\mathrm{sgs}}=\mathcal{L}_{\mathrm{gs}}+\mathcal{L}_{\mathrm{ce}},}$ with $\mathcal{L}_{\mathrm{gs}}$ defined in (2).
@@
Open-Vocabulary Semantic Gaussian Splatting. The resulting distillation method enables object localization from a pre-defined set of object classes. However, it does not support open-ended queries from users, which represents an important practical use-case. Here, we extend the capability of the distillation method from the closed-set setting to the open-vocabulary setting, leveraging open-source pre-trained text encoders, e.g., CLIP [27] and BERT [4]. These pre-trained text encoders compute text embeddings for natural-language inputs, mapping these inputs to a metric space.
@@
We pre-compute the text embeddings $\theta_{\mathrm{d}}$ of each entry in the dictionary of detected object classes using the pre-trained text encoder, without any fine-tuning. At runtime, given a natural-language prompt specified by a user, FAST-Splat computes the text embedding $\theta_{\mathrm{q}}$ associated with the prompt using the pre-trained text encoder. To identify semantically-relevant objects in the scene, we compute the relevancy score [12], which represents the pairwise softmax between the cosine-similarity values computed over the dictionary embeddings $\theta_{\mathrm{d}}$ and a set of text embeddings consisting of the query embedding $\theta_{\mathrm{q}}$ (referred to as the positive embedding) and a set of canonical embeddings (referred to as negative embeddings).
@@
Given a threshold on the relevancy score, we designate all entries whose relevancy score exceeds this specified threshold as being similar to the query. We map these entries to pixels in a rendered feature map using the probability distribution over all entries computed via the softmax function. In general, we can assign a label to each pixel by taking the maximum probability value; however, alternative techniques can be utilized, including those that provide a multi-hypothesis prediction set, e.g., conformal prediction [34].
@@
Fine-grained Language Semantics. In the real-world, identifying only the location of a related object of interest, e.g., in semantic object localization, is generally inadequate. Identifying the precise semantic class of the localized object is of even greater importance, especially for downstream applications where semantic object localization constitutes a single component of a solution pipeline. Prior semantic Gaussian Splatting methods lack this critical capability. For example, existing semantic Gaussian Splatting methods will localize a coffee machine when prompted with the query “tea,” without disambiguating the semantic identity of the localized object, which could be misleading to a user. FAST-Splat seeks to address this limitation. The novel distillation technique employed by FAST-Splat endows the algorithm with the capability to not only identify the precise semantic class of each object, but also to disambiguate the semantic identity of related objects. Given a natural-language prompt, FAST-Splat identifies the related object classes from its dictionary and subsequently, provides the location of these objects in the scene along with their specific semantic label, illustrated in LABEL:fig:teaser. We summarize the architecture in Figure 1.

### v2

@@ Abstract
We present FAST-Splat for fast, ambiguity-free semantic Gaussian Splatting, which seeks to address the main limitations of existing semantic Gaussian Splatting methods, namely: slow training and rendering speeds; high memory usage; and ambiguous semantic object localization. We take a bottom-up approach in deriving FAST-Splat, dismantling the limitations of closed-set semantic distillation to enable open-set (open-vocabulary) semantic distillation. Ultimately, this key approach enables FAST-Splat to provide precise semantic object localization results, even when prompted with ambiguous user-provided natural-language queries. Further, by exploiting the explicit form of the Gaussian Splatting scene representation to the fullest extent, FAST-Splat retains the remarkable training and rendering speeds of Gaussian Splatting. Precisely, while existing semantic Gaussian Splatting methods distill semantics into a separate neural field or utilize neural models for dimensionality reduction, FAST-Splat directly augments each Gaussian with specific semantic codes, preserving the training, rendering, and memory-usage advantages of Gaussian Splatting over neural field methods. These Gaussian-specific semantic codes, together with a hash-table, enable semantic similarity to be measured with open-vocabulary user prompts and further enable FAST-Splat to respond with unambiguous semantic object labels and $3$D masks, unlike prior methods. In experiments, we demonstrate that FAST-Splat is $6$x to $8$x faster to train, achieves between $18$x to $51$x faster rendering speeds, and requires about $6$x smaller GPU memory, compared to the best-competing semantic Gaussian Splatting methods. Further, FAST-Splat achieves relatively similar or better semantic segmentation performance compared to existing methods. After the review period, we will provide links to the project website and the codebase.
@@ 1 Introduction
Research breakthroughs in vision-language foundation models have underpinned the remarkable performance of state-of-the-art object detection and classification [24, 51], segmentation [29, 16], and image captioning [27, 44]. In general, these models learn useful multi-modal image-language representations, entirely supervised with $2$D image-text pairs, all within a shared representation space. Recent work has shown that grounding the semantic knowledge encoded by vision-language foundation models in $3$D can be useful in improving semantic segmentation and localization [13, 35] (especially in models that take in low-resolution images as input), enabling applications such as $3$D scene-editing [15, 40] and robotic manipulation [28, 36]. Although these methods demonstrate compelling object segmentation performance from open-vocabulary queries, these methods fail to disambiguate between the semantic classes of unique objects that are semantically-similar to the natural language query [13]. For example, when a user queries specifically for “tea,” existing semantic Gaussian Splatting (GSplat) methods localize the coffee machine without providing any clarifying information on the identity of the localized object, suggesting that “tea” exists in the scene, when the contrary is true. This ambiguity generally arises from the fact that many vision-language foundation methods, e.g., CLIP, utilize a bag-of-words approach in measuring the semantic similarity between image-language queries. Moreover, many semantic methods for Gaussian Splatting are notably slow, requiring a significant amount of memory and computation time for training and inference. We seek to derive a semantic GSplat method that addresses these limitations.
@@
In this work, we propose FAST-Splat, a Fast, Ambiguity-free Semantics Transfer method for Gaussian Splatting. FAST-Splat generates fine-grained semantic localization, with highly specific semantic classes even in the presence of ambiguous natural-language prompts, and as its name implies, runs much faster than existing semantic GSplat methods. FAST-Splat utilizes the vision-language foundation model CLIP [27] to embed open-set semantics into GSplat models, enabling optional initialization of relevant objects from closed-set object detectors. However, noting the limited coverage of closed-set object detectors, FAST-Splat leverages open-set object detectors and segmentation models to enable queryable semantics from free-form natural language, illustrated in Figure 1. By constructing a hash-table of identified objects, FAST-Splat enables effective semantic disambiguation, while addressing the weaknesses of closed-set methods. To enable significantly faster training and rendering speeds compared to prior work, FAST-Splat implements two key innovations: (i) single-phase semantic distillation and (ii) neural-free semantic radiance fields. FAST-Splat distills semantics into the underlying GSplat model for scene reconstruction using a single-phase training procedure, simultaneously training the geometric, visual, and semantic components of the GSplat model. In contrast, prior work, e.g., [26], utilizes a two-phase training procedure, training the geometric and visual components of the scene before introducing and optimizing the semantic parameters, negatively impacting computational performance. Moreover, unlike existing semantic GSplat methods which rely on neural-field models for semantic distillation, FAST-Splat directly augments each Gaussian with specific semantic codes, without any neural model. These key insights not only enable FAST-Splat to exploit the explicit scene-representation of Gaussian Splatting for faster training and rendering speeds, but also enable FAST-Splat to provide unambiguous semantic object labels in response to open-vocabulary user queries, e.g., in semantic object localization, upon identifying semantically-similar objects using the semantic codes associated with each Gaussian, along with the semantic hash-table.
@@
We compare FAST-Splat to existing semantic GSplat methods [26, 50] on standard benchmark datasets. FAST-Splat achieves $6$x to $8$x faster training times and $18$x to $51$x faster rendering speeds, while requiring $6$x smaller GPU memory. Meanwhile, FAST-Splat achieves competitive or better semantic object localization performance. Notably, FAST-Splat enables semantic disambiguation in object localization, even with ambiguous user-provided natural-language queries, which we illustrate in LABEL:fig:teaser. In contrast to prior work, FAST-Splat provides a clarifying semantic label to each localized object, disambiguating the semantic identity of the localized object. For example, in LABEL:fig:teaser, when prompted with the query “tea,” FAST-Splat notifies a user that it localized a coffee machine, and an apple when given the query: “fruit.” Such disambiguation can be critical to enabling interesting downstream applications, e.g., in robotics.
@@
  • We introduce FAST-Splat, a Fast, Ambiguity-Free Semantic GSplat method that enables notably faster ($6$x to $8$x faster) language-semantics grounding in $3$D scenes, distilled from $2$D vision models.
@@
  • FAST-Splat achieves superior rendering speeds for semantic Gaussian Splatting, about $18$x to $51$x faster rendering speeds compared to prior work.
@@ 2 Related Work
Grounding Language in 3D. To bridge the gap between $2$D object detection and $3$D object localization, prior work has examined grounding semantic embeddings from pre-trained image and text encoders in the $3$D world. CLIP-Fields [35] learns an implicit spatial mapping from $3$D points to CLIP image embeddings and Sentence-Bert [30] text embeddings using back-projected RGB-D points. VLMaps [11] and NLMap [4] distill vision-language features into grid-based maps. The sparsity of these maps limits the resolution of the open-vocabulary segmentation results achievable by these methods. For high-resolution object segmentation, Distilled Features Fields (DFF) [15], CLIP-NeRF [40], and LERF [13] ground CLIP [27], LSeg [16], and DINO [46] features in a NeRF representation of the scene. Moreover, similar techniques have been employed in [26, 10, 50, 52] to distill semantic features into GSplat representations. The resulting distilled radiance fields (DRFs) enable dense open-vocabulary object segmentation in $3$D and have been applied to scene-editing [15, 40] and robotic mapping and manipulation [28, 36, 49, 37, 38]. Despite their success in object segmentation, existing DRFs fail to precisely identify the semantic class of the segmented object, e.g., for a given query, these methods localize all objects related to the query without resolving the unique semantic identity of each object. This ambiguity creates a gap in scene understanding, which is often required in many downstream tasks, e.g., in robotic exploration or manipulation. In this work, we introduce a DRF with fine-grained semantics, enabling precise semantic identification of objects in a scene, disambiguating language-guided semantic object localization, and addressing existing challenges.
@@ 3 Preliminaries
Figure 1: Unlike prior semantic Gaussian Splatting methods, FAST-Splat jointly optimizes the geometric, visual, and semantic attributes of Gaussian Splatting models, achieving faster training and rendering times with effective semantics disambiguation using an efficient semantics extraction procedure.
@@ 4 Fast, Ambiguity-Free Semantic Transfer
where ${\lambda\in[0,1]}$ denotes a weighting parameter between the ${\ell_{1}}$-loss term and the differentiable structural similarity index measure (D-SSIM) loss term, and $\mathcal{I}$ and $\hat{\mathcal{I}}$ denote the ground-truth and predicted RGB image, respectively. The GSplat model is trained using the Adam optimizer [14], a gradient-based optimization method.
@@
Almost all existing semantic distillation methods rely on neural architectures for grounding language in $3$D in NeRFs [40, 15, 13] and GSplat scenes [26, 10, 50, 52]. Although effective in enabling open-set semantic object localization, neural-based approaches introduce significant computational bottlenecks into the training pipeline of NeRF and GSplat scenes. In particular, when utilized in Gaussian Splatting, neural-based semantic distillation essentially erases the computational gains, e.g., faster training times and rendering rates, afforded by Gaussian Splatting when compared to NeRFs. Moreover, prior semantic GSplat methods, e.g., [26], train the semantic attributes of a scene independently from its geometric and visual attributes, utilizing a two-phase training procedure, where the geometric and visual attributes are first optimized before the semantic attributes. This training framework significantly increases the time required for training GSplat models. We contrast this training framework with the training framework utilized by FAST-Splat, which implements a single-phase training procedure, jointly optimizing all attributes of the scene, in Figure 1.
@@
Via FAST-Splat, we seek to derive a method for open-set semantic distillation, while preserving the faster training times and rendering rates achieved by the underlying GSplat method. To achieve this goal, we revisit the fundamental insight exploited by the authors of the original GSplat work [12], namely: that only representing occupied space explicitly can lead to significant speedup in training and rendering times without any notable degradation in the visual fidelity of the representation. Specifically, Gaussian Splatting harnesses the explicit form and sparsity of an ellipsoidal-primitive-based scene representation to achieve faster training times and rendering times compared to NeRFs, the state-of-the-art at the time, which relied on implicit fields for the color and density of the scene. By learning neural encoders for dimensionality reduction or neural fields for language semantics, existing semantic distillation methods for Gaussian Splatting fail to take advantage of this insight. In contrast, in FAST-Splat, we leverage this insight to speed up training and rendering times, by eschewing neural models. In particular, we augment the attributes of each ellipsoid with a semantic parameter, associated with its semantic identity ${\beta\in\mathbb{R}^{3}}$, which is trained simultaneously with the ellipsoid’s other attributes. Note that the dimension of $\beta$ does not pose a limitation unique to this work. In fact, existing semantic GSplat methods, e.g., [26], generally use three-dimensional semantic codes to enable the utilization of the rasterization procedure employed in Gaussian Splatting.
@@
Closed-Set Semantic Gaussian Splatting. FAST-Splat supports initialization of the semantics extraction procedure using closed-set object detectors. To obtain the closed-set semantic categories, we feed each image ${I\in\mathcal{D}}$ into an object detector, e.g., YOLO [41], DETR [3], and Efficient DETR [43], to generate a list of objects present in the image. Although closed-set object detectors generally exhibit remarkable performance, these detectors occasionally fail to identify all instances of an object in an image, given that the size of their dictionary of detectable objects is often limited. To overcome this challenge, FAST-Splat utilizes an open-set detector, e.g., GroundingDINO [22], with input prompts from, e.g., image-tagging models [48] or closed-set object detectors. We examine the importance of the open-set object detector in ablations in the experiments in Section 5.
@@
At this stage, each image has a list of constituent objects and bounding boxes specifying the location of these objects in the image. However, high-fidelity semantic distillation generally requires pixel-wise semantic classes. To generate dense semantic maps, we utilize SAM-2 [29], passing in each object class and its associated bounding box. We maintain a dictionary (hash-table), representing the set of objects identified in the image. Now, we can distill the pixel-wise semantic information into the underlying GSplat model. To predict the semantic class of each ellipsoid in the scene, we apply an affine transformation $\mathcal{W}$ to the semantic attribute $\beta$ of each ellipsoid, mapping $\beta$ to an $N$-dimensional space, where $N$ denotes the size of the dictionary. Subsequently, we apply the softmax function to the resulting outputs to define a probability over the entries in the dictionary. We extend this procedure to view-based rendering. Given a camera pose for rendering, we utilize differentiable tile-based rasterization [12] to render a semantic feature map corresponding to the camera pose. We utilize the aforementioned procedure to transform the semantic feature of each pixel to a probability distribution over the entries in the hash-table. We optimize the semantic attributes using the multi-class cross-entropy loss function, given by:
@@
where $\mathcal{T}$ represents the set of all indices of pixels in the rendered feature map $\mathcal{F}_{\mathcal{I}}$ associated with image $\mathcal{I}$ (after applying the affine transformation $\mathcal{W}$), $c$ denotes the true class of the pixel, and $\gamma_{p}$ represents the weight associated with pixel $p$. We augment the set of object classes for each view with an undetected class label and assign a lower weight to pixels with this class label, i.e., pixels that were not segmented during the data pre-processing phase. This augmentation step serves as a regularization term for the stability of undetected pixels. We train all attributes of the GSplat representation, e.g., the geometric, visual, and semantic attributes and $\mathcal{W}$, simultaneously using the loss function: ${\mathcal{L}_{\mathrm{sgs}}=\mathcal{L}_{\mathrm{gs}}+\mathcal{L}_{\mathrm{ce}},}$ with $\mathcal{L}_{\mathrm{gs}}$ defined in (2).
@@
Open-Vocabulary Semantic Gaussian Splatting. The resulting distillation method enables object localization from a pre-defined set of object classes. However, it does not support open-ended queries from users, which represent an important practical use-case. Here, we extend the capability of the distillation method from the closed-set setting to the open-vocabulary setting, leveraging open-source pre-trained text encoders, e.g., CLIP [27] and BERT [5]. These pre-trained text encoders compute text embeddings for natural-language inputs, mapping these inputs to a metric space.
@@
We pre-compute the text embeddings $\theta_{\mathrm{d}}$ of each entry in the dictionary of detected object classes using the pre-trained text encoder, without any fine-tuning. At runtime, given a natural-language prompt specified by a user, FAST-Splat computes the text embedding $\theta_{\mathrm{q}}$ associated with the prompt using the pre-trained text encoder. To identify semantically-relevant objects in the scene, we compute the relevancy score [13], which represents the pairwise softmax between the cosine-similarity values computed over the dictionary embeddings $\theta_{\mathrm{d}}$ and a set of text embeddings consisting of the query embedding $\theta_{\mathrm{q}}$ (positive embedding) and a set of canonical embeddings (negative embeddings).
@@
Given a threshold on the relevancy score, we designate all entries whose relevancy score exceeds this specified threshold as being similar to the query. We map these entries to pixels in a rendered feature map using the probability distribution over all entries computed via the softmax function. In general, we can assign a label to each pixel by taking the maximum probability value; however, alternative techniques can be utilized, including multi-hypothesis prediction methods, e.g., conformal prediction [34].
@@
Fine-grained Language Semantics. In the real-world, identifying only the location of a related object of interest, e.g., in semantic object localization, is generally inadequate. Identifying the precise semantic class of the localized object is of even greater importance, especially for downstream applications where semantic object localization constitutes a single component of a solution pipeline. Prior semantic GSplat methods lack this critical capability. For example, existing semantic GSplat methods will localize a coffee machine when a user asks specifically for some “tea,” without disambiguating the semantic identity of the localized object, which could be misleading to a user. FAST-Splat seeks to address this limitation. The novel distillation technique employed by FAST-Splat endows the algorithm with the capability to not only identify the precise semantic class of each object, but also to disambiguate the semantic identity of related objects. Given a natural-language prompt, FAST-Splat identifies the related object classes from its dictionary and subsequently, provides the location of these objects in the scene along with their specific semantic label, illustrated in LABEL:fig:teaser.